A turbine operates under steady flow conditions receiving steam at the following state: pressure 1.2 MPa, temperature 188 C, enthalpy 2785 kJ/kg, velocity 33.3 m/s and elevation 3m. The steam leaves the turbine at the following state: pressure 20 kPa, enthalpy 2512 KJ/kg, velocity 100 m/s and elevation 0 m. . Heat loss to the surroundings is at a rate of 0.29 kJ/s. If the rate of steam flow through the turbine is 0.42 kg/s, what is the power output of the turbine in kW?


1
Expert's answer
2021-09-29T09:54:00-0400


Step 1

given data

At the inlet of the turbine

= 1200 kPa

= 188 C

= 2785 kJ/kg

= 33.3 m/s

= 3 m

= 0.29 kW

m ˙ = 0.42 kg/s

And at the outlet of the turbine

= 20 kPa

= 2512 kJ/kg

= 100 m/s

= 0 m

To determine

Power output of the turbine in kW

Step 2

From the steady-state equation for turbine

"m[h_i + \frac{v_1^2}{2000} + \frac{gz_i}{1000} ] + Q_r = m[h_2 + \frac{v^2_0}{2000} + gz_0] + W \\

0.42[2785 + \frac{33.3^2}{2000} + \frac{3 \times 9.81}{1000}] = 0.42[2512 + \frac{100^2}{2000} + \frac{9.81 \times 0}{1000}] + W \\

169.655 =1057.14 +W \\

W = 112.5 \; kW"

Thus, output power of turbine is 112.5 kW